David Barnes and his flancée Valerie Shah are visiting Hawail. At the Hawalian Cultural Center in Honolulu, they are told that 2 out of a group of 20 people will be randomly picked for a free lesson of a Tahitian dance. a. What is the probability that both David and Valerie get picked for the Tahitian dance lesson? (Round your answer to 4 decimal places.) b. What is the probability that Valerie gets picked before David for the Tahitian dance lesson? (Round your answer to 4 decimal places.)
Answer #1

a)P(both David and Valerie got picked)=2C2/20C2 =1/190 =0.0053

b)probability =0.5 (as either one gets selected first has equal probability)